OpenStudy (anonymous):

could someone explain finding cos(11pi/12) using Sum and difference Identities. I am a bit hung up with the initial part. I know my ref angle is 15 degrees and angle is 165 degrees. Now finding something for this I did (45-30) to use as but Im not sure that part is right. my answer of sqrt2/4(1+sqrt3) is not correct. Thanks ahead of time.

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

\[\frac{2}{3}+\frac{1}{4}=\frac{11}{12}\]

OpenStudy (misty1212):

the the addition angle formula with one angle \[\frac{2\pi}{3}\] and the other \[\frac{\pi}{4}\]

OpenStudy (anonymous):

not the answer needed. it has to use the identity \[\cos (\alpha+\beta)=\cos (\alpha)\cos(\beta) - \sin(\alpha)\sin(\beta)\] or \[\cos (\alpha+\beta)=\cos (\alpha)\cos(\beta) + \sin(\alpha)\sin(\beta)\] Im a bit hung up figuring out what \[(\alpha) and (\beta) should be though\]